find the inverse of y = (2^x) -3
Answer #1

Kindly go through the solution and let me know in case of any doubt or further clarification is needed in the comment box.
Thanks for the question :)
У: 21 - 3 g+ 3 = 2* Taking log of both side И са, (9+3)= Хаа, * х = fоu (t) f=tx)= log «)